when a number is decreased by 30% the result is 140, what is the number?
when 50% of a number is added to it the result if 180, what is the number?
thank you very much..

Answer by sumanth18(2)

when a number is decreased by 30% the result is 140, what is the number?
Lets begin by assuming the original number to be X.
Then according to the statement :

X-(30/100)*X= 140
Multiply be denominator 100 on both sides we get,
100X-30X= 14000
70X=14000
i.e X=14000/70=200
Therefore the value of X=200
Cross Verification
------------------
30% of 200 is 200*30/100=60
Decreasing 200 by 60 we get, 200-60 =140

2) when 50% of a number is added to it the result if 180
Let original number be X.
In this case we get,
X+(50/100)X=180
Solving using above way we get X=120
Cross Verification
------------------
50% of 120 is 120*50/100=60
Adding 60+120=180